OpenStudy (anonymous):

What is the slope of the line that is perpendicular to the line whose equation is 3x + y = 4?

OpenStudy (anonymous):

first put the equation in y=mx+b form

OpenStudy (anonymous):

m is your slope

OpenStudy (abmon98):

y=-3x+4 m=-3 m*m1=-1 -3*m1=-1 m1(of perpendicular)=-1/-3